# references.bib

### v1

@@ Acknowledgments
S4 Supplementary figures
Figure S1: Reconfiguring a $8\times 8$ Totimorphic lattice from a Poisson’s ratio of 00 to $-2$.
@@
S5 Supplementary videos
S5.1 Video S1a caption
Using gradient descent, the Totimorphic lattice reconfigures to attain a Poisson’s ratio of $-0.5$.
S5.2 Video S1b caption
Using gradient descent, the Totimorphic lattice reconfigures to attain a Poisson’s ratio of $0.5$.
S5.3 Video S2a caption
Compression experiment with the Totimorphic configuration with Poisson’s ratio of $-0.5$.
S5.4 Video S2b caption
Compression experiment with the Totimorphic configuration with Poisson’s ratio of $0.5$.
S5.5 Video S3a caption
Initial deployment of the Totimorphic lattice from a collapsed to a flat configuration, which is then reconfigured to focus light at a distance of $20D$, where $D$ is the side length of the lattice. On the left, a zoom-in of the lattice is shown. The origin of the lattice is also a trainable parameter, which leads to minor movement of the lattice. This could be reduced by increasing the regularisation strength (or choosing the central point in the lattice as the origin and freezing it during optimisation). On the right, the reflected light is shown. The red dot indicates the target focal point.
S5.6 Video S3b caption
Reconfiguration of the Totimorphic lattice from a focal length of $2D$ to $1D$. On the left, a zoom-in of the lattice is shown. On the right, the reflected light is shown. The red dot indicates the target focal point.

### v2

@@ 4.7 Simulation details: Totimorphic telescope
Acknowledgments
We would like to thank Derek Aranguren van Egmond, Michael Mallon, and Max Bannach for helpful and inspiring discussions, and our colleagues at ESA’s Advanced Concepts Team for their ongoing support. We further thank Leon Williams for helpful feedback on the manuscript. AT, NR, and DD acknowledge support through ESA’s young graduate trainee and fellowship programs. DD further acknowledges support through Horizon Europe’s Marie Sklodowska-Curie Actions (Project 101103062 — BASE).
Author contributions:
Initial idea is by AT and DD. Theory and model implementation is by DD, based on initial work by AT. DD, AT, NR, JG, and DI contributed to the design of the study and the performed experiments. DD and NR implemented code for the proof of concept experiments. DD wrote the first draft of the paper. DD, AT, NR, JG, and DI read and reviewed the final paper.
Competing interests:
There are no competing interests to declare.
Data and code availability:
Implementations of the models in Python are available on GitHub [8].
Supplemental Information
Supplemental figures
Figure S1: Reconfiguring a $8\times 8$ Totimorphic lattice from a Poisson’s ratio of 0 to 2-2.
@@ Acknowledgments
Supplemental movie captions
Caption for Movie S1a.
Lattice reconfiguring to attain negative Poisson’s ratio. Using gradient descent, the Totimorphic lattice reconfigures to attain a Poisson’s ratio of 0.5-0.5.
Caption for Movie S1b.
Lattice reconfiguring to attain positive Poisson’s ratio. Using gradient descent, the Totimorphic lattice reconfigures to attain a Poisson’s ratio of $0.5$.
Caption for Movie S2a.
Reconfigured lattice with negative Poisson’s ratio under load. Compression experiment with the Totimorphic configuration with Poisson’s ratio of 0.5-0.5.
Caption for Movie S2b.
Reconfigured lattice with positive Poisson’s ratio under load. Compression experiment with the Totimorphic configuration with Poisson’s ratio of $0.5$.
Caption for Movie S3a.
Deployment of a Totimorphic telescope mirror through a series of reconfigurations. Initial deployment of the Totimorphic lattice from a collapsed to a flat configuration, which is then reconfigured to focus light at a distance of $20D$, where $D$ is the side length of the lattice. On the left, a zoom-in of the lattice is shown. The origin of the lattice is also a trainable parameter, which leads to minor movement of the lattice. This could be reduced by increasing the regularisation strength (or choosing the central point in the lattice as the origin and freezing it during optimisation). On the right, the reflected light is shown. The red dot indicates the target focal point.
Caption for Movie S3b.
Adjusting telescope properties through lattice reconfiguration. Reconfiguration of the Totimorphic lattice from a focal length of $2D$ to $1D$. On the left, a zoom-in of the lattice is shown. On the right, the reflected light is shown. The red dot indicates the target focal point.